# Learning parameter dependence for Fourier-based option pricing with tensor trains

### v5

@@ V.2.1 The case of varying $\vec{\sigma}$
$d$ $\eta$ $\epsilon_{\mathrm{SVD}}$ $r_{\mathrm{TT}}$ $r_{\mathrm{MC}}+\sigma_{\text{MC}}$ $c_{\mathrm{TT}}$ $c_{\mathrm{MC}}$ $t_{\mathrm{TT}}$[sec] $t_{\mathrm{MC}}$[sec] $\chi_{\phi}$ $\chi_{\hat{v}}$ $\chi_{\tilde{V}}$
$5$ 0.9 $6.0\times 10^{-5}$ 0.0162 0.0207 $43$ $5.0\times 10^{5}$ $2.95\times 10^{-6}$ $2.84\times 10^{-2}$ $4$ $7$ $4$
$6$ 0.75 $5.0\times 10^{-5}$ 0.0231 0.0254 $75$ $6.0\times 10^{5}$ $4.30\times 10^{-6}$ $3.38\times 10^{-2}$ $5$ $7$ $5$
$7$ 0.69 $1.8\times 10^{-5}$ 0.0203 0.0236 $70$ $7.0\times 10^{5}$ $4.86\times 10^{-6}$ $3.71\times 10^{-2}$ $6$ $7$ $4$
$8$ 0.66 $2.0\times 10^{-5}$ 0.0179 0.0253 $73$ $8.0\times 10^{5}$ $5.21\times 10^{-6}$ $4.09\times 10^{-2}$ $6$ $7$ $4$
$9$ 0.61 $1.0\times 10^{-5}$ 0.0240 0.0258 $89$ $9.0\times 10^{5}$ $5.62\times 10^{-6}$ $4.51\times 10^{-2}$ $7$ $7$ $4$
$10$ auto $5.0\times 10^{-6}$ 0.0309 0.0357 $198$ $1.0\times 10^{6}$ $6.48\times 10^{-6}$ $5.11\times 10^{-2}$ $7$ $7$ $6$
$11$ auto $3.5\times 10^{-6}$ 0.0341 0.0347 $270$ $1.1\times 10^{6}$ $8.41\times 10^{-6}$ $5.04\times 10^{-2}$ $8$ $7$ $6$
@@
$d$ $\eta$ $\epsilon_{\mathrm{SVD}}$ $r_{\mathrm{TT}}$ $r_{\mathrm{MC}}+\sigma_{\text{MC}}$ $c_{\mathrm{TT}}$ $c_{\mathrm{MC}}$ $t_{\mathrm{TT}}$[sec] $t_{\mathrm{MC}}$[sec] $\chi_{\phi}$ $\chi_{\hat{v}}$ $\chi_{\tilde{V}}$
$5$ 0.9 $2.0\times 10^{-4}$ 0.0328 0.0374 $81$ $5.0\times 10^{5}$ $3.24\times 10^{-6}$ $2.90\times 10^{-2}$ $4$ $6$ $6$
$6$ 0.75 $7.0\times 10^{-5}$ 0.0265 0.0279 $114$ $6.0\times 10^{5}$ $4.32\times 10^{-6}$ $3.28\times 10^{-2}$ $5$ $6$ $6$
$7$ 0.64 $3.0\times 10^{-5}$ 0.0308 0.0355 $139$ $7.0\times 10^{5}$ $4.58\times 10^{-6}$ $3.66\times 10^{-2}$ $6$ $7$ $6$
$8$ 0.57 $1.0\times 10^{-5}$ 0.0259 0.0340 $120$ $8.0\times 10^{5}$ $5.98\times 10^{-6}$ $3.93\times 10^{-2}$ $7$ $7$ $5$
$9$ 0.53 $1.0\times 10^{-5}$ 0.0311 0.0345 $163$ $9.0\times 10^{5}$ $7.23\times 10^{-6}$ $4.54\times 10^{-2}$ $7$ $7$ $5$
$10$ 0.5 $3.0\times 10^{-6}$ 0.0315 0.0414 $171$ $1.0\times 10^{6}$ $7.99\times 10^{-6}$ $4.93\times 10^{-2}$ $7$ $8$ $5$
$11$ auto $1.0\times 10^{-7}$ 0.0379 0.0401 $125$ $1.1\times 10^{6}$ $6.27\times 10^{-6}$ $5.28\times 10^{-2}$ $9$ $14$ $4$
Table 2: The results of TT-based option pricing incorporating (a) $\vec{\sigma}$ and (b) $\vec{S}_{0}$ dependence. Here, we set the ranges $\sigma_{m}\in[0.15,0.25]$ and $S_{m,0}\in[90,120]$, and place 100 equally spaced grid points within these ranges. The $\epsilon_{\mathrm{SVD}}$ in the table refers to selected values that reduce $c_{\mathrm{TT}}$ keeping $r_{\mathrm{TT}}$ within the error range of MC-based option pricing $r_{\mathrm{MC}}+\sigma_{\text{MC}}$. The term auto in (a) and (b) indicate that $\eta_{m}$ was automatically determined for each asset based on Eqs. (18) and (19), respectively. The computational complexity and time increases when $\eta$ is set by Eq. (18) or (19). $t_{\mathrm{TT}}$ and $t_{\mathrm{MC}}$ represent the average computational time from 100 measurements, respectively.
Figure 2: The computational complexity of TT-based and MC-based option pricing, denoted as $c_{\mathrm{TT}}$ and $c_{\mathrm{MC}}$, versus the number of assets $d$. We consider including parameter dependencies on $\vec{\sigma}$ and $\vec{S_{0}}$ in TTs.
@@
V.2.2 The case of varying $\vec{S_{0}}$
TABLE 2 (b) shows the results of TT-based option pricing when we consider $\vec{S_{0}}$ dependence. TT-based option pricing demonstrates superiority over the Monte Carlo method in terms of computational complexity and time. For $d$=5 and 10, after compression using SVD and contractions, the bond dimensions between tensors related to $p_{m}$ and $z_{m+1}$ were reduced to around 10 (refer to Fig. 3(b)). The maximum bond dimension, $\chi_{\tilde{V}}$, of $\tilde{V}_{k_{1},\ldots,k_{d}}$ is 6 or less. Therefore, we saw superiority over the Monte Carlo method in terms of computation complexity and time.
Compared to varying $\sigma$, its accuracy is worse at $d=10$.
V.2.3 Randomness in learning the TTs
@@
V.2.4 Total computational time for obtaining the TTs
Lastly, let us mention the total computational time for obtaining the TTs for $\phi$ and $\hat{v}$ and $\tilde{V}_{k_{1},\ldots,k_{d}}$ through TCI and SVD. We focus on TCI because it dominates over SVD. It took about 3.4 hours for $d=11$ in the case that $\vec{\sigma}$ dependence is involved. This is sufficiently short for a practical use-case of the TT-based method mentioned exemplified in Sec. VI. We also note that, when we do not incorporate the parameter dependence into the tensor trains for $\phi$ and $\hat{v}$ as in [6], TCI takes a much longer time than the MC-based method, 7.3 seconds for $d=11$ in the case of $\vec{\sigma}$ dependence. In this setup, we set $\eta=0.48$, $\epsilon_{\mathrm{TCI}}=10^{-6}$ and all other parameters are the same as those in the default settings listed in Table 1. This means that running TCI every time the parameter varies does not lead to the time advantage of the TT-based method over the MC-based one.
VI Summary and discussion
We propose a method that employs a single TCI to learn TTs incorporating parameter dependence from functions of Eqs. (9) and (11), enabling fast option pricing in response to varying parameters. In this study, we considered scenarios with varying volatility and present stock prices as benchmarks for our proposed method. Up to $d=11$, we demonstrated superiority in both computational complexity and time. Here, we argue that, in principle, given at least comparable computational complexity, TT-based option pricing has an advantage in terms of computational time. This is because the TT-based method involves only multiplication and addition as unit operations, while the MC-based method requires more complex operations, such as random number generation and exponentiation, to be iterated $dN_{\rm path}$ times. Note that in the MC-based method, the implementation may have room for improvement to reduce computational time.
We also numerically verified that all TT-based results fell within the statistical error range of the Monte Carlo method with $10^{5}$ paths. However, it should be noted that increasing $d$ worsens accuracy, indicating a tendency for the accuracy to be worse than that of the MC-based method around $d\geq 12$.
@@ VI Summary and discussion
Finally, we discuss future research directions. If we could keep $\eta$ constant when $d$ is increased, we have the advantage of reducing the computational complexity involved in option pricing. It may be possible to keep $\eta$ constant by adjusting the value of parameter $N$, e.g. by increasing $N$ from the current value of 50 to, say, 100. It could be a promising approach to directly learning the TTOs incorporating parameter dependence using TCI, potentially improving the accuracy of option pricing. At the same time, it is important to study methods to improve the numerical stability and accuracy of TCI itself. For our method to be more practical, compressing all the input parameters including both $\vec{\sigma}$ and $\vec{S}_{0}$ into a single TN format is desired, although this study tried to incorporate either of them as the first step. Expanding the methodology to other types of option calculations, particularly the application to American options, is also an important research direction.
Acknowledgements
@@
Appendix A How to set the tolerance $\epsilon_{\mathrm{SVD}}$
Figure 4: The maximum relative error $r_{\mathrm{TT}}$ and computational complexity $c_{\mathrm{TT}}$ when applying SVD with various $\epsilon_{\mathrm{SVD}}$ values to the TTs under the parameter settings of $d=10,\vec{\sigma},\eta=\text{auto}$. The left vertical axis represents $r_{\mathrm{TT}}$, while the right vertical axis shows $c_{\mathrm{TT}}$.
Figure 4 shows the maximum relative error $r_{\mathrm{TT}}$ when applying SVD with varying $\epsilon_{\mathrm{SVD}}$ to the TT with parameters for $d=10,~{}\vec{\sigma}$, obtained through TCI. Here, we fixed the tolerance of SVD $\epsilon^{\tilde{V}}_{\mathrm{SVD}}$ to $10^{-10}$. The tolerance $\epsilon_{\mathrm{SVD}}$ for SVD was chosen to maintain $r_{\mathrm{TT}}$ smaller than $r_{\mathrm{MC}}+\sigma_{\text{MC}}$. From Fig. 4, it can be seen that $r_{\mathrm{TT}}$ increases sharply between $\epsilon_{\mathrm{SVD}}=10^{-6}$ and $10^{-5}$, suggesting that setting $\epsilon_{\text{SVD}}$ around $5.0\times 10^{-5}$ is appropriate for keeping $r_{\mathrm{TT}}$ smaller than $r_{\mathrm{MC}}+\sigma_{\text{MC}}$. By compressing the bond dimension with SVD, the computational time for the contraction of these optimized TTs also can be kept low.
From the fact that we can keep the maximum relative error $r_{\mathrm{TT}}$ small through SVD, it is suggested that the TTs obtained by TCI contain redundant information. By using SVD to get an optimal approximation in terms of the Frobenius norm, the reduced redundant information could be effectively removed. In addition, it is surprising that in the analysis with $\epsilon_{\text{SVD}}=10^{-6}$, the maximum relative error decreased compared to before compression by SVD. We consider that the error contained in TTs obtained by TCI is eliminated through SVD by chance. We expect that this phenomenon does not occur generally, and in fact, it did not occur for other asset numbers or parameters.

### v6

@@ VI.0.1 The case of varying $\vec{\sigma}$
$d$ $\eta$ $\epsilon^{\phi,\hat{v}}_{\mathrm{SVD}}$ $e_{\mathrm{TT}}$ $e_{\text{MC},10^{5}}$ $e_{\text{MC},10^{6}}$ $c_{\mathrm{TT}}$ $c_{\mathrm{MC},10^{5}}$ $t_{\mathrm{TT}}$[s] $t_{\mathrm{MC},10^{5}}$[s] $\chi_{\phi}$ $\chi_{\hat{v}}$ $\chi_{\tilde{V}}$
$5$ 0.9 $1.0\times 10^{-5}$ 0.00560 0.0213 0.00670 $14$ $5.0\times 10^{5}$ $7.77\times 10^{-7}$ $2.84\times 10^{-2}$ 12 9 2
$6$ 0.75 $1.0\times 10^{-5}$ 0.00377 0.0143 0.00456 $18$ $6.0\times 10^{5}$ $9.37\times 10^{-7}$ $3.38\times 10^{-2}$ 13 9 2
$7$ 0.69 $1.0\times 10^{-5}$ 0.00380 0.0124 0.00382 $22$ $7.0\times 10^{5}$ $9.12\times 10^{-7}$ $3.71\times 10^{-2}$ 13 8 2
$8$ 0.66 $1.5\times 10^{-5}$ 0.00547 0.0128 0.00346 $26$ $8.0\times 10^{5}$ $1.75\times 10^{-6}$ $4.09\times 10^{-2}$ 14 8 2
$9$ 0.61 $1.5\times 10^{-5}$ 0.00694 0.0100 0.00302 $30$ $9.0\times 10^{5}$ $2.47\times 10^{-6}$ $4.51\times 10^{-2}$ 13 7 2
$10$ auto $1.5\times 10^{-6}$ 0.00230 0.00916 0.00282 $34$ $1.0\times 10^{6}$ $2.00\times 10^{-6}$ $5.11\times 10^{-2}$ 59 58 2
$11$ auto $1.5\times 10^{-7}$ 0.00117 0.00847 0.00269 $38$ $1.1\times 10^{7}$ $2.16\times 10^{-6}$ $5.04\times 10^{-2}$ 73 76 2
@@
$d$ $\eta$ $\epsilon^{\phi,\hat{v}}_{\mathrm{SVD}}$ $e_{\mathrm{TT}}$ $e_{\text{MC},10^{5}}$ $e_{\text{MC},10^{6}}$ $c_{\mathrm{TT}}$ $c_{\mathrm{MC},10^{5}}$ $t_{\mathrm{TT}}$[s] $t_{\mathrm{MC},10^{5}}$[s] $\chi_{\phi}$ $\chi_{\hat{v}}$ $\chi_{\tilde{V}}$
$5$ 0.9 $1.0\times 10^{-5}$ 0.0195 0.0420 0.0133 $14$ $5.0\times 10^{5}$ $7.67\times 10^{-7}$ $3.99\times 10^{-2}$ 14 9 2
$6$ 0.75 $1.0\times 10^{-5}$ 0.0165 0.0385 0.0121 $18$ $6.0\times 10^{5}$ $9.34\times 10^{-7}$ $5.15\times 10^{-2}$ 16 9 2
$7$ 0.69 $3.0\times 10^{-5}$ 0.0103 0.0357 0.0112 $22$ $7.0\times 10^{5}$ $9.21\times 10^{-7}$ $5.37\times 10^{-2}$ $21$ $10$ $2$
$8$ 0.57 $3.5\times 10^{-5}$ 0.0125 0.0303 0.00952 $26$ $8.0\times 10^{5}$ $1.63\times 10^{-6}$ $6.18\times 10^{-2}$ $17$ $7$ $2$
$9$ 0.53 $1.0\times 10^{-5}$ 0.0155 0.0204 0.00650 $30$ $9.0\times 10^{5}$ $1.19\times 10^{-6}$ $7.95\times 10^{-2}$ $19$ $7$ $2$
$10$ 0.5 $3.0\times 10^{-5}$ 0.0172 0.0204 0.00654 $34$ $1.0\times 10^{6}$ $1.47\times 10^{-6}$ $8.77\times 10^{-2}$ $16$ $6$ $2$
$11$ auto $1.0\times 10^{-7}$ 0.0139 0.0181 0.00579 $38$ $1.1\times 10^{6}$ $1.93\times 10^{-6}$ $8.43\times 10^{-2}$ $48$ $38$ $2$
Table 2: The results of TT-based option pricing incorporating (a) $\vec{\sigma}$ and (b) $\vec{S}_{0}$ dependence. Here, we set the ranges $\sigma_{m}\in[0.15,0.25]$ and $S_{m,0}\in[90,120]$, and place 100 equally spaced grid points within these ranges. The $\epsilon^{\phi,\hat{v}}_{\mathrm{SVD}}$ in the table refers to the values adjusted so that $c_{\mathrm{TT}}$ is reduced with $e_{\mathrm{TT}}$ kept smaller than the MC error $e_{\text{MC},10^{5}}$. The term auto in (a) and (b) indicate that $\eta_{m}$ was automatically determined for each asset based on Eqs. (19) and (20), respectively. The computational complexity and time increases when $\eta$ is set by Eq. (19) or (20). $t_{\mathrm{TT}}$ and $t_{\mathrm{MC}}$ represent the average computational time from 100 measurements, respectively.
Figure 2: The computational complexity of TT-based and MC-based option pricing, denoted as $c_{\mathrm{TT}}$ and $c_{\mathrm{MC}}$ versus the number of assets $d$. We consider including parameter dependencies on $\vec{\sigma}$ and $\vec{S_{0}}$ in TTs and the number of path in MC are $10^{5}$ and $10^{6}$.
@@
VI.0.2 The case of varying $\vec{S_{0}}$
TABLE 2 (b) shows the results of TT-based option pricing when we consider $\vec{S_{0}}$ dependence. TT-based option pricing demonstrates superiority over the Monte Carlo method in terms of computational complexity and time. For $d$=5 and 10, after compression using SVD and contractions, the bond dimensions between tensors related to $p_{m}$ and $z_{m+1}$ were reduced to around 10 (refer to FIG. 3(b)). For $d=11$, when $\eta$ is determined automatically, this bond dimension increases to around 40. The maximum bond dimension, $\chi_{\tilde{V}}$, of $\tilde{V}_{k_{1},\ldots,k_{d}}$ is 2 Therefore, we saw superiority over the Monte Carlo method in terms of computation complexity and time.
Compared with the MC-based method with $10^{5}$ paths, the TT-based mathod again gives the smaller error in every value of $d$. With $10^{6}$ paths, the TT error is worse than the MC error for many values of $d$, but still comparable.
VI.0.3 Randomness in learning the TTs
@@
VI.0.4 Total computational time for obtaining the TTs
We mention the total computational time for obtaining the TTs for $\phi$ and $\hat{v}$ and $\tilde{V}_{k_{1},\ldots,k_{d}}$ through TCI and SVD. We focus on TCI because it dominates over SVD. It took about 3.4 hours for $d=11$ in the case that $\vec{\sigma}$ dependence is involved. This is sufficiently short for a practical use-case of the TT-based method mentioned exemplified in Sec. VII. We also note that, when we do not incorporate the parameter dependence into the tensor trains for $\phi$ and $\hat{v}$ as in Ref. [6], TCI takes a much longer time than the MC-based method, 7.3 seconds for $d=11$ in the case of $\vec{\sigma}$ dependence. In this setup, we set $\eta=0.48$, $\epsilon_{\mathrm{TCI}}=10^{-6}$ and all other parameters are the same as those in the default settings listed in Table 1. This means that running TCI every time the parameter varies does not lead to the time advantage of the TT-based method over the MC-based one.
VII Discussion
We propose a method that employs a single TCI to learn TTs incorporating parameter dependence from functions of Eqs. (10) and (12), enabling fast option pricing in response to varying parameters. In this study, we considered scenarios with varying volatility and present stock prices as benchmarks for our proposed method. Up to $d=11$, we demonstrated superiority in both computational complexity and time. Note that in the MC-based method, the implementation may have room for improvement to reduce computational time.
We numerically verified that all TT-based results fell within the statistical error range of the MC-based method with $10^{5}$ paths. Compared to the MC-based method with $10^{6}$ paths, the TT-based method gave the worse errors in some cases, but was still comparable. To enhance the accuracy of the TT-based method, we may reduce the tolerance of the TCI further or change the initial randomly selected grid points of TCI.
@@ VII Discussion
Finally, we discuss future research directions. If we could keep $\eta$ constant when $d$ is increased, we have the advantage of reducing the computational complexity involved in option pricing. It may be possible to keep $\eta$ constant by adjusting the value of parameter $N$, e.g. by increasing $N$ from the current value of 50 to, say, 100. It could be a promising approach to directly learning the TTOs incorporating parameter dependence using TCI, potentially improving the accuracy of option pricing. At the same time, it is important to study methods to improve the numerical stability and accuracy of TCI itself. For our method to be more practical, compressing all the input parameters including both $\vec{\sigma}$ and $\vec{S}_{0}$ into a single TN format is desired, although this study tried to incorporate either of them as the first step. Besides, to enhance the practical benefit, we should extend our methodology so that it is applicable to more advanced settings in option pricing. With respect to the pricing model, while this paper has considered the BS model, more advanced models such as the local volatility model [38] and stochastic volatility model [39, 40] are also used in practice. Applying our method to such models, where the dependencies on the parameters in the models are incorporated into TTs, is an interesting direction. Expanding the scope to the broad types of products is also desired. While this paper focused on European-type options, more complicated ones such as American and Bermudan options are also traded widely, and pricing of such options is more time-consuming. Considering whether our method can be extended for such options will be an important and interesting challenge.
VIII Acknowledgements
@@
IX Author contributions statement
R.S. conceived the project, and R.S. and H.T. wrote the code with help from K.M. and performed numerical simulations. All authors actively discussed the results and contributed to the preparation of the manuscript.
X Appendix
X.1 How to set the tolerance $\epsilon^{\phi,\hat{v}}_{\mathrm{SVD}}$
Figure 4: The maximum error $e_{\mathrm{TT}}$ when applying SVD with various $\epsilon^{\phi,\hat{v}}_{\mathrm{SVD}}$ values to the TTs under the parameter settings of $d=10,\vec{\sigma},\eta=\text{auto}$.
Figure 4 shows the maximum relative error $e_{\mathrm{TT}}$ when applying SVD with varying $\epsilon_{\mathrm{SVD}}$ to the TT with parameters for $d=10,~{}\vec{\sigma}$, obtained through TCI. Here, we fixed the tolerance of SVD $\epsilon^{\tilde{V}}_{\mathrm{SVD}}$ to $10^{-9}$. The tolerance $\epsilon^{\phi,\hat{v}}_{\mathrm{SVD}}$ for SVD was chosen to maintain $e_{\mathrm{TT}}$ smaller than $e_{\mathrm{MC},10^{5}}$. From FIG. 4, it can be seen that $e_{\mathrm{TT}}$ increases sharply between $\epsilon_{\mathrm{SVD}}=10^{-6}$ and $10^{-5}$, suggesting that setting $\epsilon_{\text{SVD}}$ around $5.0\times 10^{-6}$ is appropriate for keeping $e_{\mathrm{TT}}$ smaller than $e_{\mathrm{MC},10^{5}}$. By compressing the bond dimension with SVD, the computational time for the contraction of these optimized TTs also can be kept low.
From the fact that we can keep the maximum relative error $e_{\mathrm{TT}}$ small through SVD, it is suggested that the TTs obtained by TCI contain redundant information. By using SVD to get an optimal approximation in terms of the Frobenius norm, the reduced redundant information could be effectively removed. In addition, it is surprising that in the analysis with $\epsilon_{\text{SVD}}=10^{-6}$, the maximum relative error decreased compared to before compression by SVD. We consider that the error contained in TTs obtained by TCI is eliminated through SVD by chance. We expect that this phenomenon does not occur generally, and in fact, it did not occur for other asset numbers or parameters.